A chemist has one solution that is 25% salt and another solution that is 10% salt. How many liters of each should the chemist use to make 120 liters of a solution that is 15% salt?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

A chemist has one solution that is 25% salt and another solution that is 10% salt. How many liters of each should the chemist use to make 120 liters of a solution that is 15% salt?
--------------
Equation:
salt + salt = salt
0.25x + 0.10(120-x) = 0.15*120
----
Multiply thru by 100 to get:
25x + 10*120-10x = 15*120
15x = 5*120
x = 40 liters (amt. of 25% solution needed in the mixture)
---
120-80 liters (amt. of 10% solution needed in the mixture)
==============================================================
Cheers,
Stan H.
===========

Answer by Alan3354(69443)

Do it like this one.
--------------------
What quantity x of a 75% acid solution must be mixed with a 25% solution to produce 500 mL of a 40% solution?
---------------------------
s = 75% acid volume
t = 25% acid volume
--------------
s + t = 500
0.75s + 0.25t = 0.4*500 = 200 --> x4
3s + t = 800
1s + t = 500
-------------- Subtract
2s = 300
s = 150 ml of 75%
t = 350 ml of 25%